# Evaluation of uncertainty estimations for Gaussian process regression based machine learning interatomic potentials

### v1

@@ 2.3 Evaluating uncertainty estimations
The main objective of this study is to analyze the relationship of uncertainty estimations, such as GPR variance, and the actual prediction error. In particular, it is of interest to understand whether uncertainty estimations of ML models can serve as error estimator. A straight forward way to analyze this is to calculate uncertainty estimations and errors of a set of $n_{\text{test}}$ unseen test configurations and plot the uncertainty estimations against the errors.
To better understand, what we could potentially expect from uncertainty estimations in an ideal setting, let us assume we would have an approach that would construct an optimal predictive distribution
@@
where the true values $y(\mathbf{x})$ related to predictions $m_{opt}(\mathbf{x})$ are actually following the distribution
@@
It is then quite immediate that, under this thought experiment, the prediction error $\epsilon(\mathbf{x})$ of the model for all test samples would behave like
@@
In other words, the error of the model would be normally distributed with a mean of 00 and a standard deviation of $u_{opt}(\mathbf{x})$. In principle, this what we assume, if we consider the standard deviation of the predictive distribution to match, or to be perfectly correlated, to the error of the model. Consequently, our subsequent evaluation of the uncertainty estimations aims at testing for a given uncertainty measure $u(\mathbf{x})$, how closely the prediction errors $\epsilon$ would fulfill
@@
If the observed error distribution closely aligns with eq. (15), valid confidence intervals, which provide a range that is expected to contain the true error with a specified probability, can be derived from the uncertainties of predictions. To validate if the errors of predictions with the same predictive variance are distributed according to eq. (15), we can separate all predictions by their uncertainty values into $n_{\text{bins}}$ equidistant bins with intervals
$I_{\beta}=\left[\left(\beta-1\right)\Delta u,\,\beta\Delta u\right),\qquad\text{with}\qquad\beta=1,...,n_{\text{bins}},$ (16)
where $\Delta u$ is the bin width. We then assume that all predictions inside one bin approximately have the center of that bin
@@
as same fixed uncertainty. To evaluate whether the true error distributions align with the model’s prediction, one can calculate the bin mean and bin standard deviation of the errors $\epsilon_{\beta}$ within each bin. The mean of $\epsilon_{\beta}$ should be zero and the standard deviation should be $u_{\beta}\left(\textbf{x}\right)$, according to eq. (15). If the bin mean and bin standard deviation align with the model prediction, and the errors are also distributed according to a Gaussian, as it is predicted, also the confidence intervals, which are derived from the predictive distributions would be correct. We will refer to plots in which we plot the uncertainty estimation against the errors and mark the bin means and standard deviations of errors as correlation plots.
To illustrate a model, which perfectly aligns with eq. (15), Fig. 1 shows a correlation plot for data, where the targets were drawn from the predictive distribution of the GPR model. The distribution of errors can differ from this case, and therefore from Eq. 15, in two ways: the calculated mean of the errors is not zero and the calculated standard deviation of the errors is not $u_{\beta}\left(\textbf{x}\right)$. If the error mean differs from zero this indicates that the predictive distribution is shifted with a constant bias compared to the true distribution of targets. If the error standard deviation is smaller or larger than $u_{\beta}\left(\textbf{x}\right)$ it indicates, that the predictive distribution is either to narrow or broad, respectively.
Figure 1: Correlation plot for sinus data with synthetic GPR test targets. To generate the data a GPR model was trained with noisy data points from a sinus function. Test targets where drawn from the predictive distributions of the GPR model. Thereby data is generated for which the GPR model perfectly captures the underlying data-generating process. For all test samples the predicted GPR standard deviation is plotted against the actual error of the prediction. All predictions are separated by their standard deviation into equidistant bins. Mean and standard deviation of error distributions are shown for every bin.
@@ 2.4 Uncertainty sampling
With correlation plots, we can get a picture of different aspects that determine the quality of uncertainty estimations. As an additional study, we utilize the uncertainty estimations in an AL scheme, namely uncertainty sampling 42 to see, how effective is these estimations are in practical use. In uncertainty sampling, a small random subset of data is selected, and labels are generated for this subset. In our case, these labels are obtained via quantum chemical calculations of potential energies. This labeled data is then used to train an initial model. The uncertainty estimation of this model is used to iteratively select samples from the remaining pool of unlabeled samples for which the labels are generated and which are then added as additional training samples. See Algorithm 1 for pseudocode.
To analyze, how much uncertainty sampling with different uncertainty estimators increases data efficiency, we kept a test set separate from the data pool. In every iteration, we calculated the mean absolute error (MAE) for the test samples. Additionally, in every iteration, we calculated the maximum absolute error and the variance of the absolute errors of the test samples, to understand how consistently the model is performing across different data points. We compare adding samples via uncertainty sampling with adding samples randomly. We expect the highest increase in data efficiency, if we would have an uncertainty estimation that estimates the true error correctly. Therefore, to understand how much data efficiency could be achieved at maximum, we also iteratively added the sample with the largest absolute error in each iteration.
If, instead of adding one sample per iteration, batches of samples are added, there is a risk of adding multiple very similar samples, for all of whom the model has a high uncertainty. Although methods exist that balance informativeness, diversity, and representativeness11; 12, we do not use batch active learning approaches. Our goal to isolate the influence of the uncertainty estimation and get an accurate picture of how much additional information is extracted with different uncertainty estimations.
@@
0:  Unlabeled data pool $U$, initial sample size $n_{\text{init}}$, number of iterations $n_{\text{iter}}$
1:  Randomly select $n_{\text{init}}$ samples from $U$ and obtain their labels to form initial labeled dataset $L$
2:  $U$ = $U$$L$
@@
5:     Compute uncertainty for each sample in $U$ using model $m$
6:     Select sample $x^{*}$ from $U$ with highest uncertainty
@@ 2.5 Data
MD1749 is a benchmark dataset of long molecular dynamic trajectories, calculated with density functial theory (DFT), for 10 small molecules. The revised MD17 (rMD17) dataset50 recalculated the energies as the original values were found to be noisy. Additionally the total number of snapshots from the separate trajectory was reduced to a selection of less correlated snapshots. We selected the rMD17 dataset for our tests because it is a well-established benchmark, allowing for easy comparison and categorization of our results. We performed tests for potential energies of Benzene, as a highly symmetric molecule and Aspirin, as a more complex molecule.
The WS22 dataset51 consists of distributions of configurations obtained via Wigner sampling for 10 different organic molecules, which aims to cover an extensive configuration space. This is important for our tests to understand how uncertainty estimations can distinguish between areas of configuration spaces that are predicted accurately and those that are predicted poorly across a wide range of configurations. We performed tests for potential energies of SMA and O-HBDI.
Additionally, we considered porphyrin as a large and complex molecule with a potentially complex configuration space. The dataset consists of 100,000 snapshots of a quantum molecular dynamic trajectory calculated with Density Functional Tight Binding (DFTB). To calculate the DFTB trajectory we took the initial structure from ChemSpider52 (ID: 4086). After DFT (B3LYP/6-31G) geometry optimization, 150,000 fs ground state dynamics at 300 K were performed with the DFTB+ package53, and 100,000 frames were extracted from the end of the trajectory with a stride of 1 fs. We performed our tests for the corresponding excitation energies which we calculated with time-dependent long-range corrected DFTB (TD-LC-DFTB).
@@ 3 Results
For our experiments, we randomly select 1000 training samples and 2000 test samples from every dataset, and the remaining samples form a candidate pool. For every dataset, we use the 1000 training samples to optimize the hyperparameters. To generate the correlation plots, we calculate the uncertainty estimations and prediction errors of the respective GPR model for the candidate pool samples and plot them against each other. We separated the samples into equidistant bins by their uncertainty estimations and calculated the bin means and standard deviations of the errors. For the different datasets, we manually selected the bin width based on the range of predicted uncertainties. The exact bin widths can be found in Section S3. We compare the calculated bin means and standard deviation with the theoretical values of Eq. 15. We only show the bin mean and standard deviation, if the number of samples inside one bin is greater than 50. For better visibility, we manually adjust the axis limits of the correlation plots, therefore a few scatter points are not displayed. For all uncertainty sampling tests, we use the fixed optimized hyperparameters and the test set, as described above. From all samples except the test samples we randomly choose 200 samples as initial training samples.
3.1 rMD17 benzene
3.1.1 Correlation between uncertainty estimate and prediction error
(a) GPR with Coulomb
(b) GPR with SOAP
Figure 3: Correlation of uncertainty and error across different uncertainty estimations for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset. The samples were separated into equidistant bins by their uncertainty and the mean and standard deviation of the errors of all samples in one bin were calculated and are compared with the theoretical values from Eq. 15.
Fig. 3 shows correlation plots of different uncertainty estimators for benzene of the rMD17 dataset. In Fig. 3(a), results for GPR with Coulomb representations are presented. For the GPR standard deviation and the bootstrap uncertainty, an increasing uncertainty correlates with a broader distribution of errors, as seen by the increasing bin standard deviations. Additionally, the bin means deviate further from zero with increasing uncertainty, indicating a bias in the model’s predictions. This suggests that both uncertainty estimators can identify samples for which the model has a high bias. The correlation between higher uncertainty and more broadly distributed errors as well as a higher bias means that for predictions with higher uncertainty higher errors are more likely. Still the observed error distributions do not align with what is predicted by the GPR models, making error ranges derived from these uncertainty estimations inaccurate. The systematical bias can not be capture by the uncertainty estimations. Besides that, for the GPR standard deviation, the bin standard deviations are smaller then the theoretical values, with an increasing deviation with increasing GPR standard deviation. Therefore, the errors are more narrowly distributed than predicted, which may partly compensate for the uncaptured bias when calculating confidence intervals. For the bootstrap uncertainty, for bins with lower uncertainty, the bin standard deviations exceed the theoretical values. Therefore the errors are distributed more broadly than predicted which will result in underconfident error intervals. For bins with higher uncertainty, the bin standard deviation aligns with the theoretical values, while the bin means deviate from zero, which will result in overconfident confidence intervals. Meanwhile, the two sets uncertainty shows errors consistently distributed around zero across all bins, with nearly identical standard deviations, providing no meaningful information about the error.
In Fig. 3(b), results for GPR with SOAP representations are shown. Again for the GPR standard deviation and the bootstrap uncertainty, an increasing uncertainty correlates with broader distributions of errors, therefore higher errors become more likely with increasing uncertainty. In contrast to GPR with Coulomb matrices, the bin means for both uncertainty estimators are close to zero and do not systematically increase with increasing uncertainty. Therefore, the GPR model has almost no systematical bias in the prediction. For the GPR standard deviation also the bin standard deviations are close to the theoretical value, therefore confidence intervals for predictions derived of the uncertainty would be rather accurate. For the bootstrap aggregation, the predicted error distributions for low uncertainties are slightly too broad, while error predictions of samples with higher uncertainties become too narrow with increasing discrepancy to the theoretical value for increasing uncertainty. This results in a less significant correlation between the uncertainty and a high likelihood for high errors than observed with the GPR standard deviation. Similar as for Coulomb matrices, the error distributions for the bins of the two set uncertainty do not differ for different uncertainties. Therefore, the two sets uncertainty again does not provide any meaningful information about the error.
3.1.2 Active learning via uncertainty sampling
Fig. 4 shows uncertainty sampling results for benzene of the rMD17 dataset. In Fig. 4(a), results for GPR with Coulomb representations are shown and in Fig. 4(b) results for GPR with SOAP representations are shown. Adding samples based on maximum uncertainty estimations does not lead to any improvement in MAE, and in most cases, it also fails to improve the maximum absolute error and test variance. Only in the range of 900 to 1200 training samples for GPR with SOAP, the model with randomly added samples shows worse test variance and maximum absolute error compared to the model where samples with the highest uncertainty estimations were added. The correlation analysis shows how some of the uncertainty estimators provide relevant information about the error. E.g. when the GPR standard deviation is used as uncertainty estimator, a higher uncertainty correlates with a higher likelihood for larger errors. For GPR with SOAP, the observed error distributions actually align well with what is predicted with the GPR standard deviation. For GPR with Coulomb, there is a correlation between predictions with the highest GPR standard deviations and predictions that have the highest bias. Still, neither for GPR with Coulomb nor for GPR with SOAP uncertainty sampling with GPR standard deviation does result in higher data efficiency. In both cases, it can be seen that adding samples with the maximum absolute error improves all metrics significantly, compared to adding samples randomly. This shows that, in theory, there is a significant potential for a higher data efficiency. With the absolute error, or a potential uncertainty estimator which linearly correlates with the error, a prediction with a high error has a high uncertainty. In contrast, if the uncertainty estimator is the standard deviation of a predictive distribution, the errors of predictions with a high uncertainty follow a distribution in which high errors have a high likelihood but still also low errors are likely. This means many samples identified as maximally uncertain actually have a low error. By adding these samples, the model becomes certain about them, but it was predicting them with a low error already before. This might explain why no improved data efficiency is observed with the GPR standard deviation or bootstrap uncertainty as uncertainty estimator.
(a) GPR with Coulomb
(b) GPR with SOAP
Figure 4: Uncertainty sampling results for GPR with Coulomb and SOAP representations for benzene from the rMD17 dataset. We randomly selected 200 samples as initial training samples and 2000 samples as test samples. From the remaining pool of samples, in every iteration, the sample with the highest respective uncertainty was selected and added as an additional training sample. Additionally, we randomly added a new sample in every iteration, and we added the sample with the highest actual absolute error. During each iteration, we compute the MAE, the maximum absolute error and the variance for the test set.
3.2 Comparison for several molecules
3.2.1 Correlation between uncertainty estimate and prediction error
In Figure 5 correlation plots for the GPR standard deviation for different datasets are shown. In Figure 5(a), results for GPR with Coulomb are presented. In all cases, the bin means of bins with higher uncertainties deviate from zero, indicating a systematical bias of the respective predictions. Like it was observed for benzene, there is a correlation between increasing GPR standard deviation and increasing bias. In most cases, the bin standard deviation aligns rather well with the theoretical values. However, for SMA, it is observed that the bin standard deviations for bins with higher GPR standard deviations are smaller than the theoretical values. The deviation increases with increasing GPR standard deviation. The same can be observed to a smaller extent for O-HBDI. Therefore, the errors are more narrowly distributed than predicted. This may partly compensate for the uncaptured bias when calculating confidence intervals. Still, it can be observed that for predictions with high GPR standard deviation, the high bias dominates, which would lead to overconfident confidence intervals.
In Fig. 5(b), results for GPR with SOAP are presented. In all datasets, most of the bin means differ from zero, indicating that predictions with the GPR standard deviation of such bins have a bias. For SMA and O-HBDI, an increasing bias correlates with increasing uncertainty. For Aspirin, a clear correlation between increasing uncertainties and increasing bias cannot be observed. For SMA and Aspirin, the bin standard deviations align rather well with the theoretical values. Still, because of the uncaptured bias, resulting confidence intervals would be overconfident. For O-HBDI, the bin standard deviations for bins of low uncertainty align with the theoretical values, while for bins of high uncertainty the error distribution is much broader than predicted by the GPR model. This, combined with the uncaptured bias, would lead to highly overconfident confidence intervals.
Overall, the correlation between higher uncertainties and more broadly distributed errors, as well as a higher bias in most cases, show that there is a clear correlation between uncertainty and a higher likelihood for higher errors. Still the bin means and bin standard deviations of the error distributions do often not align with the theoretical values which are predicted by the GPR model, therefore no accurate confidence intervals can be derived from the GPR standard deviation. Specifically, predictions with a high GPR standard deviation often have a high bias, which dominates the error and is not captured by the GPR standard deviation, which would lead to highly overconfident confidence intervals for such predictions.
Here, we only compare correlation plots for the GPR standard deviation across different datasets, as it is the most widely used and discussed uncertainty estimator in the context of GPR based MLIPs. Results for two sets and bootstrap aggregation uncertainty estimations can be found in the supplementary material Section S4. In all cases, for the two sets uncertainty, the same picture as for benzene was observed: in all bins, the errors are distributed around zero with almost the same bin standard deviation. Therefore, again, the two sets uncertainty does not provide any meaningful information about the actual error. Similarly, correlation plots for the bootstrap aggregation uncertainty for Aspirin and SMA show very similar error distributions for all uncertainties, although a slight correlation between uncertainty and how much the bin means differ from zero, as well as how broad the error distributions are, can be observed. For O-HBDI, the bin standard deviations are larger than the theoretical values, with a correlation between increasing deviation and increasing uncertainties. The bin means of bins with large uncertainties also differ from zero.
3.2.2 Active learning via uncertainty sampling
In Fig. 6, the MAE for every iteration of uncertainty sampling for different datasets is shown. For all cases, beside porphyrin with Coulomb matrices, it can be observed that uncertainty sampling with the GPR standard deviation results in a worse MAE than adding samples randomly. Among all cases, porphyrin with Coulomb matrices is the only one where uncertainty sampling results in a similar MAE than adding samples randomly. In contrast to the other cases, similar to benzene with SOAP, porphyrin showed an error distribution in the correlation plot that aligns well with the theory. The two sets uncertainty performs very similar to adding samples randomly, what was expected as the correlation plots show how they do not provide any meaningful information about the error. Uncertainty sampling with the bootstrap aggregation uncertainty leads to a worse MAE compared to random sampling in all cases beside SMA with SOAP. Results for the maximum absolute test error and the test variance can be found in Section S5. The maximum absolute error of the test set improves for uncertainty sampling with GPR standard deviation in all cases, besides for aspirin. The test variance in most cases is similar to random sampling or deteriorates, only for O-HBDI a slight improvement can be observed. When training samples with the highest absolute error were added, all three metrics improved significantly for all cases, showing that in principle there is quite some potential for an improved data efficiency.
(a) GPR with Coulomb
(b) GPR with SOAP

### v2

@@ 2.3 Evaluating uncertainty estimations
The most important aspect to analyze uncertainty estimations, such as GPR standard deviation, is their relation to the actual prediction error. To better understand, what we could potentially expect from uncertainty estimations in an ideal setting, let us assume an approach that constructs an optimal predictive distribution
@@
where the true values $y(\mathbf{x})$ corresponding to predictions $m_{opt}(\mathbf{x})$ actually follow the distribution
@@
It is then quite immediate that the prediction error $\epsilon(\mathbf{x})$ of the model for all test samples would behave like
@@
In other words, the error of the model would be normally distributed with a mean of 00 and a standard deviation of $u_{opt}(\mathbf{x})$. Consequently, our subsequent evaluation of the uncertainty estimations aims at testing for a given uncertainty measure $u(\mathbf{x})$, how closely the prediction errors $\epsilon$ would fulfill
@@
The extent to which the empirical distribution of prediction errors aligns with this ideal distribution is referred to as calibration. A common method for evaluating calibration is the use of calibration curves introduced by Kuleshov et al.[37]. For every prediction within a test set, the uncertainty estimation is used to calculate a confidence interval, providing a range expected to contain the true value with a specific probability $\alpha_{\text{predicted}}$. By iterating over different $\alpha_{\text{predicted}}$ values from 0 to 1, the proportion of errors $\alpha_{\text{observed}}$ falling within the corresponding confidence intervals is calculated and plotted against $\alpha_{\text{predicted}}$. We use the implementation provided in the uncertainty toolbox[51] to generate calibration curves.
A limitation of calibration curves is that they evaluate only global calibration over the entire test set, without accounting for local calibration, e.g., differences in calibration for specific groups of predictions. To address this, we employ a method similar to the reliability diagrams proposed by Levi et al.[39]. Predictions are grouped based on their uncertainty values into $n_{\text{bins}}$ bins with intervals
$I_{\beta}=\left[\left(\beta-1\right)\Delta u,\,\beta\Delta u\right),\qquad\text{with}\qquad\beta=1,...,n_{\text{bins}}.$ (15)
Different to Levi et al., we use equidistant bins, with a fixed bin size $\Delta u$, and assume that all predictions inside one bin approximately have the center of that bin
@@
as same fixed uncertainty.
To evaluate whether the true error distributions align with the model’s predictions, we calculate the bin standard deviation of the errors $\epsilon_{\beta}$ within each bin and compare it with $u_{\beta}\left(\textbf{x}\right)$, which is the predicted standard deviation according to eq. (14). This evaluates whether the errors are as broad or narrow as predicted, but not if they are systematically shifted. In contrast to Levi et al., we therefore also calculate the bin mean of the errors $\epsilon_{\beta}$ within each bin. If the error mean differs from zero, this indicates a systematical shift of errors and suggests that the predictive distribution is shifted with a constant bias compared to the true distribution of targets. We refer to plots where uncertainty estimations are plotted against the errors, with bin means and standard deviations of errors marked, as extended reliability diagrams. To illustrate a model perfectly aligned with eq. (14), Fig. 1 shows an extended reliability diagrams for data where the targets were drawn from the predictive distribution of the GPR model.
Figure 1: Extended reliability diagram for sinus data with synthetic GPR test targets. To generate the data, a GPR model was trained with noisy data points from a sinus function. Test targets where drawn from the predictive distributions of the GPR model. Thereby data is generated for which the GPR model perfectly captures the underlying data-generating process. For all test samples, the predicted GPR standard deviation is plotted against the actual error of the prediction. All predictions are separated by their standard deviation into equidistant bins. Mean and standard deviation of error distributions are shown for every bin.
@@ 2.4 Uncertainty sampling
Beside explicitly evaluating calibration, we evaluate uncertainty estimations when utilized in an AL scheme, namely uncertainty sampling [44], to see, how effective these estimations are in practical use. In uncertainty sampling, a small random subset of a fixed dataset is selected, and labels are generated for this subset. In our case, these labels are obtained via quantum chemical calculations of potential energies. This labeled data is then used to train an initial model. The uncertainty estimation of this model is used to iteratively select samples from the remaining pool of unlabeled samples for which the labels are generated and which are then added as additional training samples. See Algorithm 1 for pseudocode.
To analyze, how much uncertainty sampling with different uncertainty measures increases data efficiency, we kept a test set separate from the data pool. In every iteration, we calculated the mean absolute error (MAE) for the test samples. Additionally, in every iteration, we calculated the maximum absolute error and the variance of the absolute errors of the test samples, to understand how consistently the model is performing across different data points. We compare adding samples via uncertainty sampling with adding samples randomly. We expect the highest increase in data efficiency, if we would have an uncertainty estimation that estimates the true error correctly. Therefore, to understand how much data efficiency could be achieved at maximum, we also iteratively added the sample with the largest absolute error in each iteration.
If, instead of adding one sample per iteration, batches of samples are added, there is a risk of adding multiple very similar samples, for all of whom the model has a high uncertainty. Although methods exist that balance informativeness, diversity, and representativeness[11, 12], we do not use batch active learning approaches. Our goal to isolate the influence of the uncertainty estimation and get an accurate picture of how much additional information is extracted with different uncertainty estimations.
@@
0:  Unlabeled data pool $U_{\text{pool}}$, initial sample size $n_{\text{init}}$, number of iterations $n_{\text{iter}}$
1:  Randomly select $n_{\text{init}}$ samples from $U_{\text{pool}}$ to get $U_{\text{init}}$ and obtain their labels to form initial labeled dataset $L$
2:  $U_{\text{pool}}$ = $U_{\text{pool}}$$U_{\text{init}}$
@@
5:     Compute uncertainty for each sample in $U_{\text{pool}}$ using model $m$
6:     Select sample $x^{*}$ from $U_{\text{pool}}$ with highest uncertainty
@@ 2.5 Data
MD17[52] is a benchmark dataset of long molecular dynamic trajectories, calculated with density functional theory (DFT), for 10 small molecules. Christensen et al.[53] recalculated the energies as the original values were found to be noisy. Additionally, the total number of snapshots from the separate trajectory was reduced to a selection of less correlated snapshots. These changes result in the revised MD17 (rMD17) dataset. We selected the rMD17 dataset for our tests because it is a well-established benchmark, allowing for easy comparison and categorization of our results. We performed tests for potential energies of Benzene, as a highly symmetric molecule and Aspirin, as a more complex molecule.
The WS22 dataset[54] consists of distributions of configurations obtained via Wigner sampling for 10 different organic molecules, which aims to cover an extensive configuration space. This is important for our tests to understand how uncertainty estimations can distinguish between areas of configuration spaces where predictions are accurate, and regions where predictions are poor across a wide range of configurations. We performed tests for potential energies of 2-(methyliminomethyl)phenol (SMA) and 4-(2-hydroxybenzylidene)-1,2-dimethyl-1H-imidazol-5(4H)-one (O-HBDI).
Additionally, we considered porphyrin as a large and complex molecule with a potentially complex configuration space. The dataset consists of 100,000 snapshots of a quantum molecular dynamic trajectory calculated with Density Functional Tight Binding (DFTB). To calculate the DFTB trajectory we took the initial structure from ChemSpider[55] (ID: 4086). After DFT (B3LYP/6-31G) geometry optimization, 150,000 fs ground state dynamics at 300 K were performed with the DFTB+ package[56], and 100,000 frames were extracted from the end of the trajectory with a stride of 1 fs. We performed our tests for the corresponding excitation energies, which we calculated with time-dependent long-range corrected DFTB (TD-LC-DFTB). Figure 2 gives the chemical structures of all molecules used in this study.
@@ 3 Results
For our experiments, we randomly select 1000 training samples and 2000 test samples from every dataset, and the remaining samples form a candidate pool. For every dataset, we use the 1000 training samples to optimize the hyperparameters. For the calibration curves and extended reliability diagrams, we calculate the uncertainty estimations and prediction errors on the candidate pool, with the respective GPR models trained on the 1000 training samples. In the extended reliability diagrams we separate the samples into equidistant bins by their uncertainty estimations and calculate the bin means and standard deviations of the errors. For the different datasets, we manually select the bin width based on the range of predicted uncertainties. The exact bin widths can be found in table S1 of the supplementary material. We compare the calculated bin means and standard deviation with the theoretical values of Eq. 14. We only show the bin mean and standard deviation, if the number of samples inside one bin is greater than 50. For better visibility, we manually adjust the axis limits of the extended reliability diagrams, therefore a few scatter points are not displayed. For all uncertainty sampling tests, we use the fixed optimized hyperparameters and the test set, as described above. For the initial model, only 200 training samples are used, randomly selected from the dataset excluding the 2,000 test samples.
3.1 Poor global calibration for ensemble uncertainties
(a) Benzene
(b) SMA
Figure 3: Calibration curves of different uncertainty measures of GPR with SOAP for rMD17 benzene and WS22 SMA. Calibration curves plot the predicted proportion of test data expected to fall within $\alpha$-prediction intervals (horizontal axis) against the observed proportion of test data that actually falls within these intervals (vertical axis), iterating over different values of $\alpha$.
In Fig. 3, calibration curves for all three uncertainty measures are shown. We exemplary show calibration curves of GPR with SOAP for rMD17 benzene and WS22 SMA. The calibration curve for the GPR standard deviation is nearly diagonal and has a low miscalibration area, indicating relatively good calibration across the dataset. In contrast, the two sets and bootstrap uncertainty measures show much poorer calibration, with significantly higher miscalibration areas. In Fig. 4 and in Figs. S3-7 of the supplementary material calibration curves for all uncertainty measures across all molecules are shown for GPR with Coulomb as well as GPR with SOAP. In nearly all cases, the GPR standard deviation demonstrates much better global calibration, with notably lower miscalibration areas compared to the other two uncertainty measures. The only exception is GPR with Coulomb for benzene. Here, the calibration curve for the GPR standard deviation shows a relatively high miscalibration area, which exceeds that of the two sets uncertainty measure.
3.2 Local calibration analysis reveals model bias
While the calibration curves indicate good global calibration for the GPR standard deviation, this does not mean that the model provides a well-calibrated uncertainty estimation for every prediction. Fig. 4 presents calibration curves together with extended reliability diagrams for GPR with Coulomb across all molecules for the GPR standard deviation. As described above, the calibration curves indicate good global calibration for all molecules except benzene. However, when predictions are grouped by their uncertainty, the extended reliability diagrams reveal poor calibration for specific groups of predictions. While the standard deviation of errors often aligns rather closely with the theoretical value, the mean value of errors can deviate strongly from zero, particularly for predictions with high uncertainties. Assuming that the errors within separate bins, for which the means and standard deviations are calculated, are distributed somewhat close to normal, this can be interpreted as a systematic shift in the error distributions. In Fig. S8 of the supplementary material, we explicitly show the error distributions within selected bins for GPR with Coulomb applied to Aspirin and O-HBDI. Indeed, these errors rather closely follow normal distributions hat are shifted relative to the predicted normal distribution. This systematic shift leads to much higher errors than predicted and indicates a systematic bias in the respective predictions. Although this bias is not captured quantitatively by the uncertainty estimation, it can still be observed that, in most cases, there is a clear correlation between increasing GPR standard deviation and increasing bias. This means the GPR standard deviation can identify predictions with a high bias.
In ML theory, the error of a model is often divided into an irreducible part resulting from the noise in the training data, as well as a bias and a variance component. Hüllermeier and Waegeman[57] describe the GPR standard deviation as a meaningful indicator of the total model uncertainty. However, if we have a model that has a bias in the form of systematical shifts of the predictive distributions in certain areas of the input space, it is not possible for the standard deviation of the predictive distribution to quantify these shifts. In general, a bias is associated with incorrect model assumptions, making it challenging to quantify the resulting error using the flawed model itself. For ensemble-based uncertainties, Heid et al.[49] discussed that a bias cannot be quantified.
Extended reliability diagrams for the GPR standard deviation of GPR with SOAP are shown in Fig. S5 of the supplementary material. In general, similar trends can be observed, although the bias is lower, and in some cases, the standard deviation of errors deviates more strongly from the theoretical value. Extended reliability diagrams for the two sets and bootstrap uncertainty of GPR wit Coulomb as well as GPR with SOAP can be found in Figs. S3-4 and S6-7 of the supplementary material. For the two sets uncertainty, it can be observed in all cases that errors are distributed with almost identical standard deviations and means across all bins. Consequently, the reliability diagrams reveal that the two set uncertainty not only shows poor global calibration but also provides no meaningful information to differentiate between predictions with different errors. A similar observation can be made for the bootstrap uncertainty in some of the cases.